# Exploring Homogeneous and Heterogeneous Consistent Label Associations for Unsupervised Visible-Infrared Person ReID

### v2

@@ Abstract
Unsupervised visible-infrared person re-identification (USL-VI-ReID) aims to retrieve pedestrian images of the same identity from different modalities without annotations. While prior work focuses on establishing cross-modality pseudo-label associations to bridge the modality-gap, they ignore maintaining the instance-level homogeneous and heterogeneous consistency in pseudo-label space, resulting in coarse associations. In response, we introduce a Modality-Unified Label Transfer (MULT) module that simultaneously accounts for both homogeneous and heterogeneous fine-grained instance-level structures, yielding high-quality cross-modality label associations. It models both homogeneous and heterogeneous affinities, leveraging them to define the inconsistency for the pseudo-labels and then minimize it, leading to pseudo-labels that maintain alignment across modalities and consistency within intra-modality structures. Additionally, a straightforward plug-and-play Online Cross-memory Label Refinement (OCLR) module is proposed to further mitigate the impact of noisy pseudo-labels while simultaneously aligning different modalities, coupled with a Modality-Invariant Representation Learning (MIRL) framework. Experiments demonstrate that our proposed method outperforms existing USL-VI-ReID methods, highlighting the superiority of our MULT in comparison to other cross-modality association methods. The code will be available.
@@ 1 Introduction
Visible-infrared person re-identification (VI-ReID) [25, 56, 44, 51, 49, 46] aims at retrieving the same person from a set of visible/infrared gallery images when given an image from another modality. It has garnered growing interest due to its practical applications in intelligent surveillance systems. Existing VI-ReID methods have achieved remarkable performance with deep neural networks [35, 19, 16, 31, 23, 28]. However, these works are based on datasets with modality-shared annotations, which are labor-intensive and time-consuming to obtain in real-world scenarios. To relieve such issues, we investigate the unsupervised solution for VI-ReID.
@@
Although there are several unsupervised single-modality ReID methods that have achieved excellent performance (e.g., Cluster-Contrast [11], ISE [55], PPLR [8]), the direct application of these methods to VI-ReID scenarios presents a formidable challenge due to the substantial modality-gap. Images of the same pedestrian from different modalities cannot be assigned the same pseudo-label using conventional clustering-based methods. Therefore, associating cross-modality pseudo-labels is necessary for unsupervised VI-ReID. Several attempts [42, 6, 5, 46] have been made to associate the cross-modality pseudo-labels. PGM [46] and MBCCM [5] adopt a global perspective to interlink cross-modality clusters. However, they overlook the complicated, fine-grained structural information at the instance level, consequently resulting in coarse associations. To utilize the instance-level information, OTLA [42] formulates the label assignment between instance and cross-modality clusters as an Optimal Transport (OT) problem. Nevertheless, it neglects the homogeneous structural consistency, leading to intra-cluster instances in one modality being dispersed across multiple clusters in another modality. Based on the above analysis, we utilize instance-level pairwise relationships to establish reliable cross-modality label associations that maintain both homogeneous and heterogeneous structural consistency (as shown in Fig.1).
@@
Specifically, we propose a Modality-Unified Label Transfer (MULT) module (Fig.2 (a)), which exploits the full potential of both homogeneous and heterogeneous instance-level structures to associate cross-modality pseudo-labels. To begin, our MULT excavates homogeneous and heterogeneous structural information by modeling affinities derived from pairwise instance relationships in feature space. These affinities are then utilized to define both homogeneous and heterogeneous inconsistency for pseudo-labels from a global perspective. Subsequently, MULT transfers the pseudo-labels guided by the calculated affinities, with the primary aim of minimizing the inconsistency terms. During the label transfer process, each instance communicates its pseudo-label information with both its intra-modality and cross-modality counterparts. Such transfer strategy leverages detailed instance-level relationships, facilitating more precise associations compared to the direct associations of clusters. Simultaneously, the pseudo-labels preserve the homogeneous structure in feature space by minimizing the homogeneous inconsistency terms. As a result, our MULT provides reliable supervision for training.
@@
To achieve modality alignment based on pseudo-labels derived from MULT, we introduce an Online Cross-memory Label Refinement (OCLR) module (Fig.2 (c)), complemented by a Modality-Invariant Representation Learning (MIRL) framework (Fig.2 (b)). Specifically, our OCLR module is a straightforward yet effective plug-and-play component designed to alleviate the impact of the inevitable noisy pseudo-labels while further reducing the modality-gap. Specifically, it learns self-consistency among memory banks. Our MIRL framework conducts both intra-modality and cross-modality contrastive learning based on memory banks. Experimental results highlight the effectiveness of our OCLR, showcasing its applicability in various cross-modality label association methods.
Our main contributions can be summarized as follows:
@@
  • We propose a MULT module that considers instance-level context structures to provide homogeneous and heterogeneous structurally consistent cross-modality pseudo-label associations for network training.
@@
  • We design a straightforward plug-and-play OCLR module for learning cross-memory self-consistency online, alleviating the impact of noisy labels while mitigating the modality-gap.
  • Experiments on SYSU-MM01 and RegDB datasets demonstrate our method outperforms existing USL-VI-ReID methods, and our MULT generates higher-quality label associations than other methods.
@@ 2.1 Unsupervised Single-Modality ReID
Unsupervised single-modality ReID aims to learn discriminative representations for unannotated person images. Recent unsupervised methods [27, 41, 26, 52, 43, 22, 53, 11, 55] follow a self-training pipeline that alternates between generating pseudo-labels and training the network. To alleviate the impact of noisy labels, MMT [14] utilize the Mean-Teacher [39] model to revise pseudo-labels online, while RLCC [54] and PPLR [8] generating offline soft pseudo-labels to refine hard labels. SPCL [15], Cluster-Contrast [11] and ISE [55] propose constructing memory banks, which store unique prototypes for unlabeled clusters.
@@ 2.2 Unsupervised VI-ReID
Unsupervised VI-ReID aims to learn modality-invariant and identity-discriminative features for cross-modality images without annotations. Existing methods [24, 47, 5, 6, 46] mainly focus on associating cross-modality pseudo-labels. H2H [24] and ADCA [47] associate heterogeneous instances from a local perspective according to pairwise similarities. To avoid biased label associations, OTLA [42] establishes an optimal transport problem to assign cross-modality pseudo-labels uniformly for instances. To model the cross-modality relationships from a global perspective, PGM [46] and MBCCM [5] construct bipartite graphs to match cross-modality clusters. Furthermore, GRU [48] proposes a CAE module to embed the information of hierarchical domains while achieving remarkable performance.
@@
Figure 2: Framework of our proposed method. Different colors indicate different modalities. Our method alternates pseudo-label generation (Modality-Unified Label Transfer (MULT (a), described in Sec.3.1)) and network training (including Modality-Invariant Representation Learning (MIRL (b), described in Sec.3.2) and Online Cross-memory Label Refinement (OCLR (c), described in Sec.3.3)).
@@ 2.3 Affinity-Based Person ReID
In the person ReID task, some works [2, 37, 36, 30, 29, 23, 49] pay attention to the detailed relationships between pairwise instances. SSFT [30] proposes a feature transfer module to facilitate the optimization of group-wise similarities for single-modality ReID. cm-SSFT [29] models both intra-modality and cross-modality affinities to generate modality-shared features from modality-specific features. CIFT [23] proposes a novel graph module to simulate the unbalanced modality distribution.
Inspired by the widely-used GCN [21] and the above affinity-based methods, we model affinities in feature space to fully exploit instance-level relationships.
@@ 3 Methodology
We employ the two stream encoder $f_{\theta}(\cdot)$ (e.g., ResNet50 [17]) to extract visible features $\left\{\mathbf{f}^{v}_{i}|i=1,2,\cdots,N^{v}\right\}$ and infrared features $\left\{\mathbf{f}^{r}_{i}|i=1,2,\cdots,N^{r}\right\}$. In the initial training stage, following [47, 5, 46], we employ the Dual-Contrastive Learning (DCL) framework [47] to facilitate intra-modality contrastive learning as our baseline.
Our proposed method is employed during the second training stage. The framework is illustrated in Fig.2. Following the well-developed unsupervised methods [15, 11, 55, 47], we alternate between pseudo-label generation and network training. During the pseudo-label generation stage, following [47, 46, 5], we first utilize DBSCAN [13] to cluster features. Two intra-modality memory banks $\tilde{\mathbf{M}}^{v}\in\mathbb{R}^{K^{v}\times d}$ and $\tilde{\mathbf{M}}^{r}\in\mathbb{R}^{K^{r}\times d}$ are initialized by the cluster centroids of their corresponding modalities, where $K^{v}$ and $K^{r}$ denotes the number of clusters in the visible modality and the infrared modality, respectively. Then the proposed Modality-Unified Label Transfer (MULT, Fig.2 (a)) establishes reliable cross-modality label associations for network training. During the training stage, we perform Modality-Invariant Representation learning (MIRL, Fig.2 (b)) to fully leverage the supervisory signals from MULT. Additionally, an Online Cross-memory Label refinement (OCLR, Fig.2 (c)) module is proposed to further alleviate the influence of noisy labels. Our entire framework includes two training modes, i.e., V-based and R-based, where V-based denotes we utilize pseudo-labels in the visible label space to guide network training, and vice versa. Following PGM [46], we alternate these two modes during training. For convenience, we only explain the V-based training mode below.
@@ 3.1 Modality-Unified Label Transfer
Motivated by the affinity-based ReID methods [30, 29, 23], Our MULT models instance-wise affinities to generate structurally consistent pseudo-labels. In MULT, every instance holds two types of pseudo-labels, i.e., the intra-modality label (labels from the label space corresponding to the instance’s modality) and the cross-modality label (labels from the label space corresponding to another modality). We use $\mathbf{y}^{e}=\{\mathbf{\tilde{y}}^{e},\mathbf{\hat{y}}^{e}\}$ to denote the pseudo-labels, where $\mathbf{\tilde{y}}^{e}$ denotes the intra-modality labels and $\mathbf{\hat{y}}^{e}$ denotes the cross-modality labels ($e=\{v,r\}$ indicates visible and infrared modality, respectively), and $\mathbf{y}^{e}_{i}$ denotes the pseudo-label of i-th instance in modality $e$. Our MULT includes two directions, i.e., V2R and R2V. We only describe the V2R MULT, which involves the transfer between visible intra-modality labels $\tilde{\mathbf{y}}^{v}$ and infrared cross-modality labels $\hat{\mathbf{y}}^{r}$.
Label Initialization. To fully utilize the relationships between instances and each cluster centroid, we use the soft probability distribution from the corresponding memory bank as initial visible intra-modality labels $\tilde{\mathbf{y}}^{v}$:
$\tilde{\mathbf{y}}^{v}\in\mathbb{R}^{N^{v}\times K^{v}},\tilde{\mathbf{y}}^{v}_{i}(0)=\mathbf{P}(\mathbf{f}_{i}^{v}|\tilde{\mathbf{M}}^{v},\tau)\in\mathbb{R}^{K^{v}},\vspace{0mm}$ (1)
${{P}_{j}(\mathbf{f}|\mathbf{M},\tau)}=\frac{\exp\left(\mathbf{f}^{\top}\mathbf{c}_{j}/\tau\right)}{\sum_{k=1}^{K}\exp\left(\mathbf{f}^{\top}\mathbf{c}_{k}/\tau\right)},\vspace{-3mm}$ (2)
$\mathbf{P}(\mathbf{f}|\mathbf{M},\tau)=[P_{1}(\mathbf{f}|\mathbf{M},\tau),\cdots,P_{K}(\mathbf{f}|\mathbf{M},\tau)],\vspace{-1mm}$ (3)
where $K$ denotes the number of prototypes in $\mathbf{M}$. Following [42, 6], we utilize Optimal Transport Label Assignment (OTLA) to initialize infrared cross-modality labels $\hat{\mathbf{y}}^{r}$:
$\hat{\mathbf{y}}^{r}(0)=OTLA(\mathbf{f}^{r},\tilde{\mathbf{M}}^{v})\in\mathbb{R}^{N^{r}\times K^{v}},\vspace{0mm}$ (4)
where $OTLA(\mathbf{f}^{r},\tilde{\mathbf{M}}^{v})$ indicates formulating the assignment of pseudo-labels in visible modality to instances in infrared modality as the OT problem, where the cost matrix for OT is calculated by $\mathbf{f}^{r}$ and $\tilde{\mathbf{M}}^{v}$.
@@
Affinity Modeling. To incorporate instance-level relationships, we model the homogeneous and heterogeneous affinities, denoted as $\mathbf{S}^{ho(e)}$ and $\mathbf{S}^{he}$. Note that $\mathbf{S}^{ho(e)}=\{\mathbf{S}^{ho(v)},\mathbf{S}^{ho(r)}\}$ denotes the affinities within visible and infrared modalities, respectively. To enhance the utilization of contextual information, we employ the Jaccard Similarity to model the homogeneous affinities:
$\mathbf{S}^{ho(e)}\in\mathbb{R}^{N^{e}\times N^{e}},\mathbf{S}^{ho(e)}_{ij}=\frac{|\mathcal{R}(\mathbf{f}^{e}_{i},\kappa)\cap\mathcal{R}(\mathbf{f}^{e}_{j},\kappa)|}{|\mathcal{R}(\mathbf{f}^{e}_{i},\kappa)\cup\mathcal{R}(\mathbf{f}^{e}_{j},\kappa)|},\vspace{-1mm}$ (5)
@@
where $N^{e}=\{N^{v},N^{r}\}$ denotes the number of instances in modality $e=\{v,r\}$. $\mathbf{S}^{ho(e)}_{ij}$ can be regarded as the affinity between $\mathbf{f}^{e}_{i}$ and $\mathbf{f}^{e}_{j}$ in modality $e$, and $\mathcal{R}(\mathbf{f}^{e}_{i},\kappa)$ is the $\kappa$-reciprocal nearest neighbors [57] of $\mathbf{f}^{e}_{i}$.
Due to the substantial modality discrepancy, directly modeling the heterogeneous affinity by the Jaccard Similarities between cross-modality features as Eq.5 to model the heterogeneous affinity will lead to the problem of modality-bias: some instances are naturally closer to another modality so they hold a higher degree of total affinity to that modality. To address the modality-bias, we hope all instances within one modality have an equal total affinity to another modality. Thus we model the heterogeneous affinites by solving the Optimal Transport (OT) plan according to the transport cost between cross-modality features, which can be formulated as:
@@
Inconsistency Formulation. We leverage the above affinities to define the inconsistency terms for pseudo-labels. A minor degree of structural inconsistency implies that the higher the affinity between two instances, the smaller the distance between their pseudo-labels. Consequently, we define the product of the Euclidean distance between pseudo-labels of pairwise instances and their affinity as their structural inconsistency. From a global perspective, for all instances in both modalities, the homogenous inconsistency can be formulated as follows:
$\mathcal{I}^{e}_{ho}({\mathbf{y}}^{e})=\sum_{i=1}^{N^{e}}\sum_{j=1}^{N^{e}}\mathbf{S}^{ho(e)}_{ij}\|{\mathbf{y}}_{i}^{e}-{\mathbf{y}}_{j}^{e}\|^{2}_{2}.\vspace{-1.2mm}$ (7)
Here $\mathbf{y}^{e}=\{{\tilde{\mathbf{y}}}^{v},{\hat{\mathbf{y}}}^{r}\}$ denotes the pseudo-labels involved in the process of V2R MULT. The heterogeneous inconsistency can be formulated in a similar manner:

### v4

@@ Abstract
Unsupervised visible-infrared person re-identification (USL-VI-ReID) endeavors to retrieve pedestrian images of the same identity from different modalities without annotations. While prior work focuses on establishing cross-modality pseudo-label associations to bridge the modality-gap, they ignore maintaining the instance-level homogeneous and heterogeneous consistency between the feature space and the pseudo-label space, resulting in coarse associations. In response, we introduce a Modality-Unified Label Transfer (MULT) module that simultaneously accounts for both homogeneous and heterogeneous fine-grained instance-level structures, yielding high-quality cross-modality label associations. It models both homogeneous and heterogeneous affinities, leveraging them to quantify the inconsistency between the pseudo-label space and the feature space, subsequently minimizing it. The proposed MULT ensures that the generated pseudo-labels maintain alignment across modalities while upholding structural consistency within intra-modality. Additionally, a straightforward plug-and-play Online Cross-memory Label Refinement (OCLR) module is proposed to further mitigate the side effects of noisy pseudo-labels while simultaneously aligning different modalities, coupled with an Alternative Modality-Invariant Representation Learning (AMIRL) framework. Experiments demonstrate that our proposed method outperforms existing state-of-the-art USL-VI-ReID methods, highlighting the superiority of our MULT in comparison to other cross-modality association methods. Code is available at https://github.com/FranklinLingfeng/code_for_MULT.
Keywords:
Homogeneous and heterogeneous consistency Label associations Unsupervised visible-infrared person re-identification
journal: International Journal of Computer Vision
@@ 1 Introduction
Visible-infrared person re-identification (VI-ReID) Lin et al (2022); Zhang et al (2021b); Wu et al (2017); Ye et al (2021b, 2020); Wu and Ye (2023) aims at retrieving the same person from a set of visible/infrared gallery images when given an image from another modality. It has garnered growing interest due to its practical applications in intelligent surveillance systems. Existing VI-ReID methods have achieved remarkable performance with deep neural networks Pu et al (2020); Jia et al (2020); Hao et al (2019); Mao et al (2017); Li et al (2022b); Liu et al (2022). However, these works are based on datasets with modality-shared annotations, which are labor-intensive and time-consuming to obtain in real-world scenarios. To relieve such issues, we investigate the unsupervised solution for VI-ReID.
Although there are several unsupervised single-modality ReID methods that have achieved excellent performance ($e.g.$, Cluster-Contrast Dai et al (2023), ISE Zhang et al (2022b), PPLR Cho et al (2022)), the direct application of these methods to VI-ReID scenarios presents a formidable challenge due to the substantial modality-gap. Images of the same pedestrian from different modalities cannot be assigned the same pseudo-label using conventional clustering-based methods. Therefore, associating cross-modality pseudo-labels is necessary for unsupervised VI-ReID. Several attempts Wang et al (2022); Cheng et al (2023b, a); Wu and Ye (2023) have been made to associate the cross-modality pseudo-labels. PGM Wu and Ye (2023) and MBCCM Cheng et al (2023a) adopt graph matching to interlink cross-modality clusters from the global perspective. However, they overlook the complicated, fine-grained structural information at the instance level, consequently resulting in coarse associations. To utilize the instance-level information, OTLA Wang et al (2022) formulates the label assignment between instance and cross-modality clusters as an Optimal Transport (OT) problem. Nevertheless, it neglects the homogeneous structural consistency, leading to a large amount of intra-cluster instances in one modality being dispersed across multiple clusters in another modality. Based on the above analysis, we utilize instance-level pairwise relationships to establish reliable cross-modality label associations that maintain both homogeneous and heterogeneous structural consistency (as shown in Fig.1).
@@
Specifically, we propose a Modality-Unified Label Transfer (MULT) module (Fig.2 (a)), which exploits the full potential of both homogeneous and heterogeneous instance-level structures to associate cross-modality pseudo-labels. To begin, our MULT excavates homogeneous and heterogeneous structural information by modeling affinities derived from pairwise instance relationships in feature space. These affinities are then utilized to define both homogeneous and heterogeneous inconsistency between the pseudo-label space and the feature space from a global perspective. Subsequently, MULT transfers the pseudo-labels guided by the calculated affinities, with the primary aim of minimizing the inconsistency terms. During the label transfer process, each instance communicates its pseudo-label information with both its intra-modality and cross-modality counterparts. Such transfer strategy leverages detailed instance-level relationships, facilitating more precise associations compared to the direct associations of clusters. Simultaneously, the pseudo-labels preserve the homogeneous structure in feature space by minimizing the homogeneous inconsistency terms. Furthermore, the transferred soft pseudo-labels involve information between instances and multiple intra-modality and cross-modality clusters and thus are more suitable for mining discriminative features compared to hard labels. Extensive experiments indicate that our MULT provides more suitable supervision signals for training compared to other cross-association methods.
@@
To achieve modality alignment based on pseudo-labels derived from MULT, we introduce an Online Cross-memory Label Refinement (OCLR) module (Fig.2 (c)), complemented by an Alternative Modality-Invariant Representation Learning (AMIRL) framework (Fig.2 (b)). Specifically, our OCLR module is a straightforward yet effective plug-and-play component designed to alleviate the impact of the inevitable noisy pseudo-labels while further reducing the modality-gap. Specifically, it learns self-consistency among the predictions from multi-memory prototypes. Our AMIRL framework conducts contrastive learning based on both intra-modality and cross-modality memory banks. Two auxiliary memory banks are constructed to collaboratively learn the intra-modality structure with the initial two intra-modality memory banks, which play a role in mutually correcting each other. Furthermore, an alternative training scheme is proposed to multigate the influence of the inconsistency between the pseudo-labels from MULT in different directions. Experimental results highlight the effectiveness of our OCLR, showcasing its applicability in various cross-modality label association methods. Our main contributions can be summarized as follows:
@@
  • We propose a MULT module that considers instance-level context structures to provide homogeneous and heterogeneous consistent cross-modality pseudo-label associations for network training. The generated pseudo-labels exhibit cross-modality alignment while containing rich intra-modality information.
@@
  • We design a straightforward plug-and-play OCLR module for learning cross-memory self-consistency online, coupled with an AMIRL framework that fully exploits the supervision signals from MULT. The OCLR module effectively alleviates the side-effects of the noisy labels while mitigating the modality-gap.
  • Extensive experimental results demonstrate the effectiveness of the proposed method, showing higher-quality label associations across-modality instances, and superior recognition performances to state-of-the-art USL-VI-ReID methods on SYSU-MM01 and RegDB datasets.
@@ 2 Related Work
2.1 Supervised VI-ReID
Supervised VI-ReID pays attention to retrieving pedestrian images from both visible and infrared cameras. Existing methods mainly focus on aligning two modalities at the image-level and feature-level. Some methods Mao et al (2017); Dai et al (2018); Choi et al (2020) bridge the modality-gap by transferring images from one modality to the other based on Generative Adversarial Networks (GANs). However, these methods face the issue of high computational cost and the inevitable noise in generated images, thus degenerating the model performance. CA Ye et al (2021a) proposes a simple random channel argumentation for visible images to bridge the gap between cross-modality images. For aligning modalities at feature-level, some methods design novel network structures based on two-stream CNN with deep metric learning Liu et al (2022); Wu et al (2021); Ye et al (2020); Zhang and Wang (2023); Wu et al (2023) to excavate modality-invariant features. DDAG Ye et al (2020) mines both intra-modality part-level and cross-modality graph-level contextual information through a dynamic dual-attention aggregation learning method. MPANet Wu et al (2021) proposes a modality alleviation module and a pattern alignment module to extract part-level discriminative features. PartMix Kim et al (2023) synthesizes positive and negative by mixing the part descriptors across modalities to enhance part-level feature learning. However, the aforementioned methods rely heavily on human-annotated cross-modality associations, which require expensive labor costs and are not always available in real scenarios. Thus we focus on learning modality-invariant features without human annotations.
@@ 2.2 Unsupervised Single-Modality ReID
Unsupervised single-modality ReID aims to learn discriminative representations for unannotated person images. Recent unsupervised methods Lin et al (2020); Wang and Zhang (2020); Lin et al (2019); Yu et al (2019); Wang et al (2021); Li et al (2022a); Zhang et al (2023); Dai et al (2023); Zhang et al (2022b) follow a self-training pipeline that alternates between generating pseudo-labels and training the network. The mainstream methods for unsupervised ReID include pseudo-label refinement-based methods and memory bank-based methods. To alleviate the impact of noisy labels, MMT Ge et al (2020a) utilizes the Mean-Teacher Tarvainen and Valpola (2017) model to revise pseudo-labels online. RLCC Zhang et al (2021a) propagates pseudo-labels between different iterations through cluster consensus and generates temporally consistent pseudo-labels. PPLR Cho et al (2022) proposes refining the pseudo-labels of global features by ensembling the predictions of part features based on the global-part agreement score. Memory bank-based methods establish memories capable of managing multi-prototypes and facilitate contrastive learning during training. SPCL Ge et al (2020b) constructs a hybrid memory that maintains both instance-level and cluster-level prototypes with a self-paced learning strategy. Cluster-Contrast Dai et al (2023) and ISE Zhang et al (2022b) store a unique prototype for each cluster to preserve the updating consistency. DCMIP Zou et al (2023) manages discrepant cluster memories and a multi-instance memory to excavate multifaceted information within clusters.
@@ 2.3 Unsupervised VI-ReID
Unsupervised VI-ReID aims to learn modality-invariant and identity-discriminative features for cross-modality images without annotations. Existing methods Liang et al (2021); Yang et al (2022); Cheng et al (2023a, b); Wu and Ye (2023) mainly focus on associating cross-modality pseudo-labels. H2H Liang et al (2021) proposes an ISML loss aimed at enhancing alignment between reliable cross-modality instances. ADCA Yang et al (2022) identifies highly associated cross-modality features and aggregates their corresponding memory prototypes according to instance-level pairwise similarities. To avoid biased label associations, OTLA Wang et al (2022) establishes an optimal transport problem to assign cross-modality pseudo-labels uniformly for instances. To model the cluster-level cross-modality relationships from a global perspective, PGM Wu and Ye (2023) and MBCCM Cheng et al (2023a) construct weighted bipartite graphs to associate cross-modality clusters and assign shared pseudo-labels for heterogeneous instances. CCLNet Chen et al (2023) leverages the powerful semantic information from CLIP to provide richer supervision signals by incorporating a prompt learning stage. GRU Yang et al (2023) proposes a CAE module to embed the information of hierarchical domain memories while achieving remarkable performance.
@@
Figure 2: Framework of our proposed method. Different colors indicate different modalities. Our method alternates pseudo-label generation (Modality-Unified Label Transfer (MULT (a), described in Sec.3.1)) and network training (including Alternative Modality-Invariant Representation Learning (AMIRL (b), described in Sec.3.2) and Online Cross-memory Label Refinement (OCLR (c), described in Sec.3.3)). MULT provides homogeneous and heterogeneous consistent pseudo-labels as supervision signals. During training, AMIRL leverages memory banks to perform contrastive learning with an alternative scheme and OCLR utilizes predictions from different memories to alleviate the effect of the noisy labels.
@@ 2.4 Affinity-Based Person ReID
In the person ReID task, some works Chen et al (2018); Shen et al (2018b, a); Luo et al (2019); Lu et al (2020); Li et al (2022b); Ye et al (2020) pay attention to the detailed relationships between pairwise instances. SSFT Luo et al (2019) proposes a feature transfer module to facilitate the optimization of group-wise similarities for single-modality ReID. cm-SSFT Lu et al (2020) models both intra-modality and cross-modality affinities to generate modality-shared features from modality-specific features. CIFT Li et al (2022b) proposes a novel graph module to simulate the unbalanced modality distribution.
Inspired by the widely-used GCN Kipf and Welling (2016) and the above affinity-based methods, we model affinities in feature space to fully exploit instance-level relationships. Then we aim to obtain pseudo-labels that preserve instance-level contextual information.
@@ 3 Methodology
We employ the two stream encoder $f_{\theta}(\cdot)$ ($e.g.$, ResNet50 He et al (2016)) as backbone to extract visible features $\left\{\mathbf{f}^{v}_{i}|i=1,2,\cdots,N^{v}\right\}$ and infrared features $\left\{\mathbf{f}^{r}_{i}|i=1,2,\cdots,N^{r}\right\}$. In the initial training stage, following Yang et al (2022); Cheng et al (2023a); Wu and Ye (2023), we employ the Dual-Contrastive Learning (DCL) framework Yang et al (2022) to facilitate intra-modality contrastive learning as our baseline.
Our proposed method is employed during the second training stage. The framework is illustrated in Fig.2. Following the well-developed unsupervised methods Ge et al (2020b); Dai et al (2023); Zhang et al (2022b); Yang et al (2022), we alternate between pseudo-label generation and network training. During the pseudo-label generation stage, following Yang et al (2022); Wu and Ye (2023); Cheng et al (2023a), we first utilize DBSCAN Ester et al (1996) to cluster features. Two intra-modality memory banks $\tilde{\mathbf{M}}^{v}\in\mathbb{R}^{K^{v}\times d}$ and $\tilde{\mathbf{M}}^{r}\in\mathbb{R}^{K^{r}\times d}$ are initialized by the cluster centroids of their corresponding modalities, where $K^{v}$ and $K^{r}$ denotes the number of clusters in the visible modality and the infrared modality, respectively. Then the proposed Modality-Unified Label Transfer (MULT, Fig.2 (a)) establishes reliable cross-modality label associations and generates soft modality-unified pseudo-labels for network training. During the training stage, we propose an Alternative Modality-Invariant Representation learning (AMIRL, Fig.2 (b)) framework to fully leverage the supervisory signals from MULT through contrastive learning, coupled with an alternative scheme. Additionally, an Online Cross-memory Label refinement (OCLR, Fig.2 (c)) module is proposed to further alleviate the influence of noisy labels while reducing the modality discrepancy. The entire framework includes two training modes, $i.e.$, V-based and R-based, where V-based denotes we utilize pseudo-labels in the visible label space to guide network training, and vice versa.
@@ 3.1 Modality-Unified Label Transfer
Motivated by the affinity-based ReID methods Luo et al (2019); Lu et al (2020); Li et al (2022b), Our MULT models instance-wise affinities to generate structurally consistent pseudo-labels. In MULT, every instance holds two types of pseudo-labels, $i.e.$, the intra-modality label (labels from the label space corresponding to the instance’s modality) and the cross-modality label (labels from the label space corresponding to another modality). Formally, We use $\mathbf{y}^{e}=\{\tilde{\mathbf{y}}^{e},\hat{\mathbf{y}}^{e}\}$ to denote the pseudo-labels, where $\tilde{\mathbf{y}}^{e}$ denotes the intra-modality labels and $\hat{\mathbf{y}}^{e}$ denotes the cross-modality labels. $e=\{v,r\}$ indicates visible and infrared modality, respectively, and $\mathbf{y}^{e}_{i}$ denotes the pseudo-label of $i$-th instance in modality $e$. ($e.g.$, $\tilde{\mathbf{y}}^{v}$ denotes pseudo-labels for visible instances in visible label space and $\hat{\mathbf{y}}^{v}$ denotes pseudo-labels for visible instances in infrared label space).
Our MULT includes two directions, $i.e.$, V2R and R2V. The V2R MULT involves the transfer between visible intra-modality labels $\tilde{\mathbf{y}}^{v}$ and infrared cross-modality labels $\hat{\mathbf{y}}^{r}$ within the visible label space, and verse vice. For convenience, we only describe the V2R MULT in detail.
@@
Affinity Modeling. To incorporate instance-level relationships, we model the homogeneous and heterogeneous affinities, denoted as $\mathbf{S}^{ho(e)}$ and $\mathbf{S}^{he}$. $\mathbf{S}^{ho(e)}=\{\mathbf{S}^{ho(v)},\mathbf{S}^{ho(r)}\}$ denotes the affinities within visible and infrared modalities, respectively. To enhance the consistency between the homogeneous affinities and the clustering results, we employ the Jaccard Similarity Zhong et al (2017) for clustering to model the homogeneous affinities:
$\mathbf{S}^{ho(v)}_{ij}=\frac{|\mathcal{R}(\mathbf{f}^{v}_{i},\kappa)\cap\mathcal{R}(\mathbf{f}^{v}_{j},\kappa)|}{|\mathcal{R}(\mathbf{f}^{v}_{i},\kappa)\cup\mathcal{R}(\mathbf{f}^{v}_{j},\kappa)|},\mathbf{S}^{ho(v)}\in\mathbb{R}^{N^{v}\times N^{v}},$ (1)
$\mathbf{S}^{ho(r)}_{ij}=\frac{|\mathcal{R}(\mathbf{f}^{r}_{i},\kappa)\cap\mathcal{R}(\mathbf{f}^{r}_{j},\kappa)|}{|\mathcal{R}(\mathbf{f}^{r}_{i},\kappa)\cup\mathcal{R}(\mathbf{f}^{r}_{j},\kappa)|},\mathbf{S}^{ho(r)}\in\mathbb{R}^{N^{r}\times N^{r}},$ (2)
@@
where $\mathcal{R}(\mathbf{f}_{i},\kappa)$ is the $\kappa$-reciprocal nearest neighbors Zhong et al (2017) of $\mathbf{f}_{i}$, $i.e.$, $\mathcal{R}(\mathbf{f}_{i},\kappa)=\{\mathbf{g_{i}}|(\mathbf{g_{i}}\in\mathbb{N}(\mathbf{f}_{i},\kappa))\wedge(\mathbf{f}_{i}\in\mathbb{N}(\mathbf{g_{i}},\kappa))\}$, and $\mathbb{N}(\mathbf{f}_{i},\kappa))$ denotes the $\kappa$-nearest neighbors of the probe $\mathbf{f}_{i}$, ||$\cdot$|| ⋅ | is the cardinality of a set. $\mathbf{S}^{ho}_{ij}$ can be regarded as the affinity between homogeneous features $\mathbf{f}_{i}$ and $\mathbf{f}_{j}$.
Due to the substantial modality discrepancy, it is inappropriate to directly model the heterogeneous affinity by the Jaccard Similarities between cross-modality features as Eq.1 and Eq.2. It should take into account both the cross-modality instance-level relationships and the misalignment between distributions of the two modalities. Thus affinity modeling can be conceptualized as a transition process between instances from two different distributions, which can be formulated as an optimal transport problem. Each instance in the dataset should be treated equally, meaning the marginal distributions of instances from both modalities follow uniform distributions. Thus we model the heterogeneous affinites by solving the Optimal Transport (OT) plan according to the transport cost between cross-modality features, which can be formulated as:
@@
Inconsistency Formulation. We leverage the above affinities to define the inconsistency terms for pseudo-labels. A minor degree of structural inconsistency implies that the higher the affinity between two instances, the smaller the distance between their pseudo-labels. Consequently, we define the product of the Euclidean distance between pseudo-labels of pairwise instances and their affinity as their pairwise structural inconsistency. From a global perspective, for all instances in both modalities, the homogenous inconsistency can be formulated as follows:
$\mathcal{I}^{v}_{ho}(\tilde{\mathbf{y}}^{v})=\sum_{i=1}^{N^{v}}\sum_{j=1}^{N^{v}}\mathbf{S}^{ho(v)}_{ij}\|{\tilde{\mathbf{y}}}_{i}^{v}-{\tilde{\mathbf{y}}}_{j}^{v}\|^{2}_{2},$ (4)
$\mathcal{I}^{r}_{ho}(\hat{\mathbf{y}}^{r})=\sum_{i=1}^{N^{r}}\sum_{j=1}^{N^{r}}\mathbf{S}^{ho(r)}_{ij}\|{\hat{\mathbf{y}}}_{i}^{r}-{\hat{\mathbf{y}}}_{j}^{r}\|^{2}_{2}.$ (5)
The above inconsistency term can be regarded as a weighted sum of pairwise distances between the soft pseudo-labels. The higher the affinities between two instances in feature space, the more their pseudo-label distance contributes to the computation of the inconsistency. The heterogeneous inconsistency can be formulated in a similar manner: